The sum of seven consecutive integers is 980. How many of them are prime?
Answer by checkley77(12844)

LET X, X+1, X+2, X+3, X+4, X+5, X+6 BE THE 7 CONSECUTIVE INTEGERS.
X+X=1+X+2+X+3+X+4+X+5+X+6=980
7X+21=980
7X=980-21
7X=959
X=959/7
X=137 ANS. FOR THE FIRST INTEGER. PRIME
137+1=138 ANS.
137+2=139 ANS. PRIME.
137+3=140 ANS.
137+4=141 ANS.
137+5=142 ANS.
137+6=143 ANS.
PROOF:
137+138+139+140+141+142+143=980
980=980